# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v1

@@ 3 Sequential Bayesian Experimental Design for LLMs
    Construct updated joint, $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, using the new history (Section˜3.3) and return to Step 1 (unless a termination criterion like maximum number of questions has been achieved).
Because BED-LLM uses the prior–likelihood pairing, our belief state on $\theta$ is simply given by $p_{f}(\theta;h_{t})$ at any given point in the process; our underlying aim is that the uncertainty of this distribution contracts as quickly as possible.
In the remainder of this section, we provide details on these key individual steps. We note that some of the choices made are closely linked to our joint model construction; we will later provide extensive discussion on why this model construction was chosen and possible alternatives in Section˜4.
3.1 Candidate Question Generation
The first step of each BED-LLM iteration is to generate a finite pool of candidate questions, $\mathcal{X}^{\text{cand}}$, from which our chosen question will be selected. To ensure we ultimately choose good questions, we must generate a candidate set $\mathcal{X}^{\text{cand}}$ that is rich and diverse enough to include sufficiently informative questions in the first place.
We consider two approaches for generating candidate questions:
  • Unconstrained generation. Given the accumulated history, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t+1};h_{t})$ with appropriate prompting.
  • Conditional generation. The LLM is given both the accumulated history and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t+1};[h_{t},\Theta^{\text{cand}}])$. Specifically, the LLM is prompted to propose questions whose answers would “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions independently with a relatively high temperature (we use $T=1.3$ in our experiments), in order to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks over-fitting to the sampled set of hypotheses. In practice, we found that conditional generation works well for discrete, easily divisible hypothesis spaces such as those found in the 20–Questions setting, but tends to overfit in the case of more complex, overlapping hypotheses such as persona identification. We can avoid this overfitting by generating two hypothesis sets, one for generating questions and one for estimating EIG. However, this adds significant computational cost, leading us to prefer the unconstrained generation approach in such cases.
We restrict our questions to be multiple-choice questions. This is because having a small space of single-token responses $y$ (e.g., yes/no, A/B/C/D) ensures that $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ provides a meaningful uncertainty over outcomes and that we can reliably calculate/estimate its entropy. We discuss this restriction at length in Section˜4.2, where we will see that limitations in LLM uncertainty quantification over unrestricted spaces generally require us to use a restricted space for $y$ (as done here) or $\theta$ (which can be tackled by instead using the data–estimation pairing).
@@ 3.2 Estimating EIG for each question
We can estimate the EIG for a given question $x_{t+1}$ using the following Rao-Blackwellized estimator of Equation 3, independently introduced by Gal et al, (2017) and Rainforth, (2017):
@@
$\begin{split}\text{EIG}_{\theta}(x_{t+1};h_{t})\approx&\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t+1}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\log p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\\ &-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]),\end{split}$ (5)
where $\hat{p}(y_{t+1};[h_{t},x_{t+1}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ and $\theta^{(n)}\sim p_{f}(\theta;h_{t})$ (see Section˜3.3). While this estimator is biased for finite NN, [^3] In general it is not possible to derive unbiased estimates for the EIG without using MLMC approaches (Goda et al,, 2022) that significantly increase the cost-normalized variance. it is still a singly intractable estimator that converges as $N\to\infty$ at the standard Monte Carlo rate $O(1/\sqrt{N})$ (Rainforth et al,, 2018). Note that the samples $\theta^{(n)}$ do not need to be independent for this to hold, provided that they can be shown to satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley, (2013)).
@@
Note that when constructing this estimator, one should directly utilize the LLM’s logits (whenever this is possible) to calculate the $p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ terms. By the Rao-Blackwell theorem (Rao et al,, 1945), this will always produce lower variance estimators than purely sample–based estimators, like those employed in Hu et al, (2024) and Kobalczyk et al, (2025).
3.2.1 EIG cannot be well approximated using predictive entropy
Previous attempts to apply BED to the LLM setting with a prior–likelihood based model have all, either explicitly or implicitly, assumed that the expected likelihood entropy, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$, is constant over designs (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025); for BED-LLM this would equate to ignoring the first line of the estimator in Equation˜5. Under this assumption, the EIG breaks down to simply the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t})}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$. Therefore, approaches that make this assumption are no longer information–based, and are instead making decisions simply based on predictive uncertainty.
This is problematic as, in practice, this expected conditional uncertainty will vary between questions. For example, if $\theta=$“dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not. In general, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$ measures how easily the question can be answered once $\theta$ is known, and intuitively this will not be constant across questions for the vast majority of problems.
This variability in the expected entropy of the likelihood can be crucial in determining good questions and omitting it provides no meaningful computational saving (as the same LLM evaluations are used for the top and bottom lines of Equation˜5, doing the full estimate of the EIG requires no additional LLM calls to be made). We, therefore, strongly advocate against making this assumption: including the expected entropy of the likelihood is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. This recommendation will be strongly backed up by our experiments, where we will see that approximating the EIG with the marginal predictive entropy significantly degrades performance of the sequential BED approach. A further example of how predictive entropy can be detrimental is given in Figure˜1.
One might instead be tempted to replace $\hat{p}(y_{t+1};[h_{t},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$ in the EIG estimation, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$.
@@ 3.3 Prior Construction and Belief Updating
The Savage axioms (Savage,, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily over–collapsed to a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f. Figure˜8).
Question 1 Question:
Which ice cream flavor feels like the best match for this user?
Predictive Entropy: Very High
EIG: 0
Question 2 Question:
Which film genre does the user most prefer?
Choose one option: A. Action B. Sci-Fi C. Comedy D. Horror
Predictive Entropy: High
EIG: High
@@
A natural tractable alternative is to simply derive our beliefs through in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t})$, noting that this has been shown to behave differently to Bayesian updating in the LLM context (Falck et al,, 2024; Kossen et al,, 2024). Unfortunately, we find that the ability of current LLMs to fully incorporate the information from $h_{t}$, even when using large state-of-the-art LLM variants such as GPT-4o (OpenAI,, 2024) or Mistral Large (Mistral AI Team,, 2024), is surprisingly lacking. In particular, we find that all LLMs we evaluate regularly sample hypotheses $\theta$ that are incompatible with previous observations in $h_{t}$, especially as the length of the history grows. We also, again, find problems with over–collapsing, with LLMs often jumping to conclusions about the value of $\theta$ without appropriate evidence for this certainty in $h_{t}$.
The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al, 2024b, ; Zhang et al, 2025b, ; Zhang et al,, 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al,, 2022; Kossen et al,, 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
@@
Addressing these shortfalls is critical to the success of BED-LLM. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. We refer to our derived distribution as $p_{f}(\theta;h_{t})$ and its key component is to filter the generated hypotheses according to whether they are compatible with the history $h_{t}$. We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t}$ (using $p_{\mathrm{LLM}}(y;[\theta,x])$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous step t1t\!-\!1 which remain consistent with the most recent question and observation are retained in the hypothesis set for step $t$ without regeneration.
We also make some further variations from $p_{\mathrm{LLM}}(\theta;h_{t})$ to try and combat the over–collapsing issue and encourage diversity (noting that our initial sampling acts like a proposal in a rejection sampler and invalid samples should still be removed by the filtering). First, we sample at a high temperature ($T=1.3$). Second, rather than generating candidates independently, we prompt the LLM to generate batches of candidates using a prompt that requests a “diverse and representative set of candidates”. We can sample multiple batches, with the previous batches in context to force the LLM to generate different samples. Finally, we assume a uniform distribution over the $\theta$ that make it through the rejection, such that we remove any duplicated samples and do not make any direct use of the LLM’s logits on $\theta$. We note that the precise prompting strategy used in $p_{\mathrm{LLM}}(\theta;h_{t})$ can also be quite important in achieving effective performance, with our prompts set up to try and maximize the diversity in the generated $\theta$ samples. Details on our exact setup for $p_{f}(\theta;h_{t})$ are in Appendix˜B.
Under the aforementioned setup for $p_{f}(\theta;h_{t})$, the original prior $p(\theta)$ only affects the EIG at the first step as it has no direct bearing on $p_{f}(\theta;h_{t})$ for $t>0$. We, therefore, define our original prior to simply be equal to $p_{f}(\theta;h_{t})$ with $h_{t}=\emptyset$, such that no filtering is performed. We note though that in some cases it may be preferable to specify some customized $p(\theta)$, potentially then adjusting the filtering process and/or weighting the accepted samples in a non–uniform manner to ensure this prior information is not lost as more data is gathered.
4 On the Specification of $p(\theta,y_{t+1};h_{t},x_{t+1})$, and its Implications
As we described in Section˜3, adapting BED to the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t+1};h_{t},x_{t+1})$. As we discuss in this section, this choice is critical to the success of BED-LLM and impacts how we model the current uncertainty about $\theta$; how easily and accurately we can estimate EIG; and how aligned the model’s actual information gain under a question is predicted by the EIG, which ultimately governs whether the questions we choose are actually good at reducing uncertainty about $\theta$.
@@ 4 On the Specification of $p(\theta,y_{t+1};h_{t},x_{t+1})$, and its Implications
As discussed at the start of Section˜3, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t})p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}, [^4] One could in principle also consider prompting the LLM to generate $\theta$ and $y$ jointly as well, but we will omit these from our discussions on the basis that this will generally be impractical from an estimation perspective, while LLMs’ autoregressive nature means that this should be equivalent to one of the aforementioned joints, up to effects in prompting strategy and API handling etc. where each of the $p$ terms either corresponds directly to the LLM’s distribution over outputs given the provided context and suitable prompting (i.e. the corresponding $p_{\mathrm{LLM}}(\cdot;\cdot)$), some valid distribution derived from this, or an externally specified distribution. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second is analogous to a one-sample martingale posterior that directly samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning (Fong et al,, 2023; Falck et al,, 2024).
In our outlined BED-LLM approach (c.f. Section˜3), we adopted a prior–likelihood pairing. Below, we will expand on some reasons to prefer this formulation and discuss certain settings where the data–estimation might be preferable.
4.1 Modeling flexibility
@@
The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t})$, whereas the prior is only implicitly defined in the data–estimation pairing. In Section˜3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t})$, wherein we made various adjustments to $p_{\mathrm{LLM}}(\theta;h_{t})$ to encourage diversity and better incorporate information from the history.
On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. Most notably, in some applications like scientific discovery, we may have access to external simulators that can be run to predict our data, with the LLM used instead primarily to automate how we draw conclusions from such hypothetical data (Ghareeb et al,, 2025).
In the cases considered in this paper, we do not have such external, validated simulators, and the flexibility in prior construction is more important for our needs.
4.2 Faithfulness of conditional distributions
@@
In Section˜4.1, we considered flexibility in specifying the marginal distributions $p(\theta;h_{t})$ and $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ and $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, in practice, these will typically be more difficult and expensive to implement. This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t+1}$ respectively), rather than just needing us to set up one marginal distribution that can then be repeatedly sampled from. Second, to construct estimators for Equations (2) and (3), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations). As such, the conditionals need to be explicit distributions, or at least distributions where the probability can be cheaply estimated (noting that such estimation will be nested and thus more expensive than in conventional settings (Rainforth et al,, 2018)), so they are more difficult to define through the output of some algorithmic procedure, especially if the number of possible realizations is large.
@@
When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution. Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from. If we accept the LLM’s direct predictive distribution for $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$. In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply (noting the nesting means we need to repeatedly carry out this estimation for different $\theta$ and $y$ samples respectively).
@@
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al,, 2019; Paninski,, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al,, 2022; Desai & Durrett,, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For example, suppose $\theta$ represents the space of Python code a user wants to generate and $y$ is the answer to a simple yes/no question about that code’s desired behavior. Here, the LLM’s entropy over yes vs. no when given the code will be a much more meaningful uncertainty measure than its entropy over code when told the answer to the question.
For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead. We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜A.
4.3 Extracting the belief state

### v2

@@ 3 Sequential Bayesian Experimental Design for LLMs
    Construct updated joint (Section˜3.3): $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$, using the new history and return to Step 1 (unless a termination criterion has been achieved).
Note our belief state on $\theta$ after the $t$-th turn is simply given by $p_{f}(\theta;h_{t})$.
3.1 Generating Candidate Questions
As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\text{cand}}$, then select the best question from these. We consider two specific approaches: 1) Unconstrained generation. Given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting. 2) Conditional generation. The LLM is given both $h_{t-1}$ and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\text{cand}}])$. Specifically, the LLM is prompted to propose questions that “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks overfitting to $\Theta^{\text{cand}}$. In practice, we find it is effective for discrete spaces (§6.1), but less so for spaces with complex, overlapping hypotheses (§6.2). We restrict questions to multiple-choice format to allow $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ to produce well-calibrated probabilities (see §4).
@@ 3.2 Estimating EIG for each question
To estimate the EIG based on Equation 2 for a given question $x_{t}$, we derive the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
@@
$\begin{split}\text{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
@@
where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ and $\theta^{(n)}\!\sim\!p_{f}(\theta;h_{t-1})$, see §3.3. This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017). Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)). When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ terms using the LLM’s logits whenever possible. By the Rao-Blackwell theorem, this always produces lower variance than purely sample–based estimators (Rao et al., 1945), like those employed in Hu et al. (2024) and Kobalczyk et al. (2025).
Avoiding deterministic likelihood assumptions
Previous attempts to apply information criteria to choosing queries in LLMs have generally assumed responses are deterministic given $(\theta,x_{t})$ (Cooper et al., 2025; Kobalczyk et al., 2025; Hu et al., 2024; Mazzaccara et al., 2024; Piriyakulkij et al., 2023). Under this assumption, the EIG simplifies to the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$.
This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$. For example, if $\theta=$“Dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not. In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known. Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. We provide an illustrative example of this in Section˜A.2.1 and empirical evidence for this in §6. Given that approximating the EIG with marginal predictive entropy does not provide meaningful computational savings (as it does not reduce the required number of LLM calls), we advise against making such deterministic likelihood assumptions.
@@ 3.3 Prior Construction and Belief Updating
The Savage axioms (Savage, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f. Figure˜5).
@@
A natural tractable alternative is to derive our beliefs through LLM in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t-1})$, noting that this has been shown to behave differently to Bayesian updating (Falck et al., 2024; Kossen et al., 2024). However, we find that even state-of-the-art LLMs such as GPT-4o (OpenAI, 2024) often fail to incorporate history faithfully; they regularly sample hypotheses incompatible with past observations and exhibit premature overconfidence, with both issues becoming more pronounced as $h_{t-1}$ grows. We discuss reasons for these shortfalls in Section˜A.3.
@@
To avoid these shortfalls, we instead propose an approach that balances tractability and faithfulness. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. Our derived distribution, which we refer to as $p_{f}(\theta;h_{t-1})$, differs from $p_{\mathrm{LLM}}(\theta;h_{t-1})$ in two key ways. First, we filter the generated hypotheses according to whether they are compatible with the history $h_{t-1}$. We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])\forall i=1:t-1$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous turn which remain consistent with the most recent question and observation are retained in the hypothesis set without regeneration. Second, we make a number of modifications to promote diversity. Rather than generate candidates independently, we prompt the LLM to generate batches of candidates using a prompt encouraging diversity. After filtering these candidates as above and removing duplicates, we then impose a uniform distribution. Details of our exact setup for $p_{f}(\theta;h_{t-1})$ are given in Appendix˜E.
4 On the Specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its Implications
@@
As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t};h_{t-1},x_{t})$. In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t-1})p\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second has analogies to a marginal-posterior approach (Fong et al., 2023; Falck et al., 2024) in that it that samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning. In our outlined BED-LLM approach, we adopted a prior–likelihood pairing. Below, we justify this decision and also discuss certain settings where the data–estimation setup might be preferable instead.
@@
Modeling flexibility   The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t-1})$, whereas the prior is only implicitly defined in the data–estimation pairing. In §3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t-1})$. On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t};[h_{t-1},x_{t}]\right)$, which could, for example, be useful when we have access to external data simulators.
@@
Faithfulness of conditional distributions   While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t};\left[\theta,x_{t}\right]\right)$ and $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, in practice, these will typically be more difficult and expensive to implement. This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t}$ respectively), rather than just needing us to set up a single marginal distribution. Second, to construct estimators for Equations (1) and (2), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations). As such, the conditionals need to be explicit distributions, or at least ones where the probability can be cheaply estimated, so they are more difficult to define through the output of some algorithmic procedure, especially in large spaces.
@@
When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution. Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from. If we accept the LLM’s direct predictive distribution for $p\left(y_{t};\left[\theta,x_{t}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t-1},x_{t+1},y_{t+1}])]$. In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply.
@@
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al., 2019; Paninski, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al., 2022; Desai & Durrett, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead. We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜B.